Asset C3PO has a depreciable base of $18.15 million and a service life of 10 years. What would the accumulated depreciation be at the end of year five under the sum-of-the-years' digits method? (Do not round intermediate calculations.)

Multiple Choice

  • $4.95 million.

  • $9.07 million.

  • $13.20 million.

  • None of these answer choices are correct.

2.Cutter Enterprises purchased equipment for $99,000 on January 1, 2018. The equipment is expected to have a five-year life and a residual value of $6,900.

Using the sum-of-the-years'-digits method, depreciation for 2019 and book value at December 31, 2019, would be:

Multiple Choice

  • $26,400 and $32,700 respectively.

  • $24,560 and $36,840 respectively.

  • $24,560 and $43,740 respectively.

  • $26,400 and $39,600 respectively.

3.

Cutter Enterprises purchased equipment for $42,000 on January 1, 2018. The equipment is expected to have a five-year life and a residual value of $6,000.

Using the double-declining balance method, depreciation for 2019 would be:

Multiple Choice

  • $16,800.

  • $8,640.

  • $10,080.

  • None of these answer choices are correct.

    3.Cutter Enterprises purchased equipment for $42,000 on January 1, 2018. The equipment is expected to have a five-year life and a residual value of $6,000.

    Using the double-declining balance method, depreciation for 2019 would be:

    Multiple Choice

  • $16,800.

  • $8,640.

  • $10,080.

  • None of these answer choices are correct.
Answer #1

1. Depreciation under sum of years digit method

Depreciation base = 18.15 million

Year Depreciation base Remaining life of the asset Depreciation fraction Depreciation expenses Book value
1 18,150,000 10 10/55 3,300,000 14,850,000
2 18,150,000 9 9/55 2,970,000 11,880,000
3 18,150,000 8 8/55 2,640,000 9,240,000
4 18,150,000 7 7/55 2,310,000 6,930,000
5 18,150,000 6 6/55 1,980,000 4,950,000

Accumulated depreciation after 5th year is $13.20 million (3.30+2.97+2.64+2.31+1.98)

Depreciation fraction=10+9+8+7+6+5-4+3+2+1=55

2. Depreciation base = cost of thes asset - salvage value

=99,000-6,900= $92,100.

Useful life =5 years

Year Depreciation base Remaining life Depreciation fraction Depreciation expenses Book value
1 92,100 5 5/15 30,700 61,400
2 92,100 4 4/15 24,560 36,840

Depreciation fraction = 6+4+3+2+1=15

Depreciation for 2019 is $24,560 and book value is $36,480.

3. Double Declining balance method

Double Declining balance formula = 2× cost of the asset× depreciation value

Here it is (1/5)×100 = 20%

20%×2=40%

Depreciation base=42,000.

Depreciation for 2018 = 42,000×40%= 16,800

Book value at the end of the year =42,000-16,800= 25,200.

Depreciation for 2019 =25,200×40% = $10,080.

Not : in the problem, question 3 is repeated twice.

____×____

All the best

Any doubt comment below

Please support us